# Rethinking Table Instruction Tuning

### v3

@@ 4.3 Hindsight Analysis
Figure 3: Performance scores (y-axis) with respect to the number of epochs (x-axis) and learning rates. In Figure 3(a), we aggregate the performance scores for all the datasets listed in Table 5.
@@
In hindsight, we want to validate that our selected hyperparameters indeed work the best. Therefore, we run the experiments on the same training set with the learning rate ranging from 1.0e-7 to 1.0e-5, and the number of epochs from one to six. Figure 3 reports part of the results, and Appendix E reports the complete results and provide further discussion.
As shown in Figure 3(a), in the table understanding tasks, the learning rate of 1.0e-6 and 5.0e-7 yields the best overall performance, which coincides with our findings in Section 3. In addition, the model achieves its best aggregated performance around two to three epochs for both learning rate.
@@
For MMLU, both 1.0e-6 and 5.0e-7 maintain their performance, sometimes even slightly better than the original LLaMA 3.1 8B Instruct model. As revealed in Table 9, the performance boost is more pronounced on STEM category. We hypothesize that this is because table-related tasks typically involve data analysis that requires math reasoning, which belongs to the STEM category. Therefore, training on table-related tasks would lead to better STEM performance. This also explains the performance boost for MMLUPro in Table 8.
@@
Generally, the trends we observe here resemble the trends we have observed in Section 3. A learning rate that is too large or too small would lead to suboptimal performance on table understanding tasks, and fine-tuning the model with one or two epochs would result in a competitive model without the risk of sacrificing its general capabilities. Moreover, we demonstrate here that with preliminary experiments, we can find a set of good or even the best hyperparameters to train the final model. Therefore, we encourage researchers to prioritize hyperparameter selection and conduct preliminary experiments when developing their models.
@@ 6 Conclusion
Due to the space constraint, we provide further analysis across LLMs in Section D.4, including Llama 2 7B Instruct (Touvron et al., 2023), QWen 2.5 7B Instruct (Bai et al., 2023), Mistral v0.3 7B Instruct (Jiang et al., 2023), and Phi 3 small 8K Instruct (7B) (Abdin et al., 2024), and analysis in terms of LoRA and QLoRA in Section D.5. We provide further analysis regarding how the data features affect the model’s performance degradation on general benchmarks in Section D.7. However, due to the scope of one study, we cannot exhaust every possible foundational model and every possible dataset.
@@ A.1 Better TAMA Models with Limited Data
Notably, in Table 10, TAMA-QWen3 achieves the best overall performance of 33.9, surpassing QWen-3-8B (32.9) and TableGPT-2 (30.0).
@@ B.1 GPU Details
Table 11 shows the generation hyperparameters for table LLMs.
@@ B.3 Batch Sizes
We conduct preliminary experiments to evaluate the impact of batch size on model performance, testing batch sizes of 4, 8, 16, 32, and 64. These experiments are performed on a total of 600 training instances from FeTaQA, HiTab, and TabFact (200 from each dataset, respectively as detailed in Section 3). We train the Llama 3.1 8B Instruct model for three epochs with a learning rate of 1e-6.
@@
Based on these findings, we select a batch size of 16 for our experiments in Sections 3 and 4.
@@ B.6 Details of Evaluation
We provide the performance breakdown in terms of the category for MMLU in Table 12.
@@ D.1 Model Selection
Currently, there are two kinds of model selections for table instruction tuning, instruction-tuning the base version of the model, as seen in works like TableLLaMA(Zhang et al., 2024a) and TableBenchLLM(Wu et al., 2024), or continuing instruction-tuning an already instruction-tuned version, as done with TableLLM(Zhang et al., 2024b) as listed in Table 1.
As the end user may come up with their own set of instructions, we expect table instruction-tuned models to possess a strong general instruction-following ability. Imparting general instruction-following ability through table instruction-tuning to the base model is challenging, as there is a lack of diversity in the table instruction-tuning data. For instance, TableLLaMA employs six specific instruction templates across two million data points, which pales in comparison to the diverse instruction datasets in broader instruction tuning efforts such as those by Chung et al. (2024), which include 1,836 tasks, each with a set of instruction templates. As shown in Figure 4(c), when tuning the base version of the LLaMA 3.1 8B model on instruction pairs on FeTaQA, HiTab, and TabFact, the instruction following ability of the model does not improve significantly. Moreover, with a large learning rate such as 1.0e-5, the model’s instruction following ability drops significantly when there is more training data coming in.
@@
In terms of specific table understanding tasks, tuning LLaMA 3.1 8B Instruct model yields better performance than its base version on TabFact (73.10 in Figure 2(a) v.s. 71.10 in Figure 4(a)) under the same experimental setup. Therefore, we select the LLaMA 3.1 8B Instruct model as our starting model.
@@ D.2 Effects of Epochs
Figure 5 illustrates the relationship between the performance of LLaMA 3.1 8B Instruct model and the number of epochs when we fine-tune the model on the 1,500 instruction pairs at a learning rate of 1.0e-6. The model demonstrates a decent performance on these table tasks within just one or two epochs. In the meantime, the model mostly preserves its performance on MMLU and IFEval, indicating that its general capabilities are not compromised too much while acquiring table reasoning ability. Beyond this point, there is no significant performance improvement, suggesting that extending training for more epochs yields diminishing returns or may even lead to overfitting. Therefore, we choose to train our model for two epochs in Section 4 instead of the commonly adopted six epochs by existing table LLMs as seen in Table 1.
@@ D.3 Effects of Multi-Task
In Figure 6, we present the heatmap of model performance when fine-tuning the LLaMA 3.1 8B Instruct model on a single dataset (one of the datasets among FeTaQA, HiTab, and TabFact). We fine-tune the model for two epochs at a learning rate of 1.0e-6 with 500 instruction pairs, and then test it against the six datasets. Additionally, Figures 7 and 8 in Section D.6 present heatmaps across varying learning rates (from 1.0e-7 to 1.0e-5) and number of epochs (from one to six).
@@
There are synergy effects on these tasks. The model achieves better performance when trained on the instruction pairs combined from all three datasets, compared to being trained on each of them separately. For instance, the accuracy on HiTab increases to 66.29, compared to 64.84 when trained only on HiTab as shown in Figure 7.
There are inter-connections between different tasks. In Figure 6, we note that fine-tuning solely on HiTab leads to a performance of 67.80 on TabFact, and fine-tuning solely on TabFact leads to a performance of 55.62 on HiTab, demonstrating a transfer of learned capabilities between these two tasks. However, this relationship is not universal as training on HiTab yields poor performance on FeTaQA, indicating that the overlap between certain tasks may be limited.
Based on these observations, we choose to fine-tune our model on a diverse range of tasks and datasets in Section 4.
@@ D.4 Hyperparameter Exploration Across Models
We train each model on 500 examples from HiTab, FeTaQA, and TabFact (1,500 examples total) to explore the effects of the learning rate. Table 13 presents our results.
@@
Table 4 lists the learning rate we would suggest for practitioners to use if they would fine-tune the LLMs on table-specific tasks.
@@
We further experiment with various training sizes for each model to observe its impact on performance. Table 14 reports the results for Llama 2 7B, QWen 2.5, Mistral v0.3, and Phi 3 8K models at one of the learning rates we select based on our results in Table 13.
@@
In addition, we find that given the same number of training instances, Llama 3.1 8B Instruct achieves better performance than Llama 2 7B Instruct. For instance, when trained with the same 1,500 examples at the learning rate of 1.0e-6, Llama 3.1 8B Instruct yields 73.10 on TabFact (Section 3) while Llama 2 7B Instruct only yields 53.80 (Table 14). Therefore, models with stronger general capabilities require less tuning data in our fine-tuning process.
@@ D.5 Hyperparameter Exploration for LoRA and QLoRA
We replicate the experiments we conduct in Section D.4, and here we present our results in two aspects, the learning rate and the number of examples.
@@
Table 15 presents the results. We find that there is still a “breakdown point” where further increasing the learning rate causes a sharp decline in overall performance for both LoRA and QLoRA. However, such “breakdown point” for LoRA and QLoRA (around 5.0e-5) is larger than the full parameter tuning (usually around 1.0e-6). When the learning rate does not surpass such a “breakdown point”, both methods demonstrate competitive in-domain performance on table tasks.
@@
Table 16 presents the results. Similar to what we have found for full parameter fine-tuning, both LoRA and QLoRA show diminishing returns as the number of training examples increases. While performance improves with more examples, the rate of improvement slows beyond 600 examples for LoRA. For QLoRA, the rate of improvement slows beyond 90 examples. We find that with 1,500 examples, QLoRA and LoRA perform similarly on the in-domain table tasks, and on FeTaQA, QLoRA even outperforms LoRA by 1 point. This suggests that practitioners may leverage such parameter-efficient fine-tuning methods like QLoRA in practice, especially when they have limited table data.
@@ D.6 Individual Task’s Influence on Model Performance
Figures 7 and 8 present heatmaps across varying learning rates (from 1.0e-7 to 1.0e-5) and epochs (from one to six). We can see that the patterns coincide with what we have discussed in Section 3, that a learning rate that is too large such as 1.0e-5 or too small such as 1.0e-7 leads to suboptimal table understanding ability, and the large learning rate also compromises the model’s general capabilities. Moreover, we do not observe significant performance gain when we fine-tune the model for more epochs. Across these hyperparameters, we can observe the inter-connections between tasks such as HiTab and TabFact, as training solely on one often leads to good performance on the other. But this is not universally true, as tasks such as FeTaQA and FEVEROUS seem to not have strong inter-connections.
In addition, we observe that the learning rate works the best for an individual task does not necessarily work the best for other tasks. For instance, in Figures 7 and 8, the learning rate of 5.0e-6 yields the best performance for FeTaQA, but is suboptimal for HiTab and TabFact. This highlights that when multiple tasks are involved in the training process, researchers need to consider beyond a single task to decide their hyperparameters.
@@ D.7 Trade-off Analysis for Data Properties
Table 17 presents the results. We find that the performance degradation is most significant on TabFact. Interestingly, despite TabFact having intermediate numeric density and table-to-question token ratios, it still shows the fastest performance decline.
@@ Appendix E Hindsight Analysis
Figure 9: Performance scores (y-axis) with respect to the number of epochs (x-axis) and learning rates. In Figure 9(a), we aggregate the performance scores for all the datasets listed in Table 5.
Figure 9 provides the complete results of the model performance versus the learning rate and the number of epochs.
Apart from what we have discussed in Section 4.3, we find that on S1, the learning rate of 5.0e-7 yields a consistent good accuracy scores (around 64 to 65) across all the epochs, while 1.0e-6 maintains a good accuracy score (around 64 to 65) for the first two epochs, but starting from the third epoch, it experiences a performance decline (from 64.93 to 52.56).
@@ Appendix F Model Prediction Examples
Table 18 provides an example for table LLMs’ generation on IFEval dataset. Tables 19 and 20 provide two examples for table LLM’s generation on Table-Syn dataset. Apart from the limited out-of-domain table reasoning ability, we find that existing table LLMs also exhibit limited instruction-following capabilities, and often struggle with consistently returning answers in specified formats, such as JSON. Such a limitation poses challenges in the practical use cases, where the end-users may request specific output formats to extract answers from the model’s predictions.

### v4

@@ 4.3 Hindsight Analysis
Figure 3: Performance scores (y-axis) with respect to the number of epochs (x-axis) and learning rates. In Figure˜3(a), we aggregate the performance scores for all the datasets listed in Table˜6.
@@
In hindsight, we want to validate that our selected hyperparameters indeed work the best. Therefore, we run the experiments on the same training set with the learning rate ranging from 1.0e-7 to 1.0e-5, and the number of epochs from one to six. Figure˜3 reports part of the results, and Appendix˜E reports the complete results and provide further discussion.
As shown in Figure˜3(a), in the table understanding tasks, the learning rate of 1.0e-6 and 5.0e-7 yields the best overall performance, which coincides with our findings in Section˜3. In addition, the model achieves its best aggregated performance around two to three epochs for both learning rate.
@@
For MMLU, both 1.0e-6 and 5.0e-7 maintain their performance, sometimes even slightly better than the original LLaMA 3.1 8B Instruct model. As revealed in Table˜10, the performance boost is more pronounced on STEM category. We hypothesize that this is because table-related tasks typically involve data analysis that requires math reasoning, which belongs to the STEM category. Therefore, training on table-related tasks would lead to better STEM performance. This also explains the performance boost for MMLUPro in Table˜9.
@@
Generally, the trends we observe here resemble the trends we have observed in Section˜3. A learning rate that is too large or too small would lead to suboptimal performance on table understanding tasks, and fine-tuning the model with one or two epochs would result in a competitive model without the risk of sacrificing its general capabilities. Moreover, we demonstrate here that with preliminary experiments, we can find a set of good or even the best hyperparameters to train the final model. Therefore, we encourage researchers to prioritize hyperparameter selection and conduct preliminary experiments when developing their models.
@@ 6 Conclusion
Due to the space constraint, we provide further analysis across LLMs in Section˜D.4, including Llama 2 7B Instruct (Touvron et al., 2023), QWen 2.5 7B Instruct (Bai et al., 2023), Mistral v0.3 7B Instruct (Jiang et al., 2023), and Phi 3 small 8K Instruct (7B) (Abdin et al., 2024), and analysis in terms of LoRA and QLoRA in Section˜D.5. We provide further analysis regarding how the data features affect the model’s performance degradation on general benchmarks in Section˜D.7. However, due to the scope of one study, we cannot exhaust every possible foundational model and every possible dataset.
@@ A.1 Better TAMA Models with Limited Data
Notably, in Table˜11, TAMA-QWen3 achieves the best overall performance of 33.9, surpassing QWen-3-8B (32.9) and TableGPT-2 (30.0).
@@ B.1 GPU Details
Table˜12 shows the generation hyperparameters for table LLMs.
@@ B.3 Batch Sizes
We conduct preliminary experiments to evaluate the impact of batch size on model performance, testing batch sizes of 4, 8, 16, 32, and 64. These experiments are performed on a total of 600 training instances from FeTaQA, HiTab, and TabFact (200 from each dataset, respectively as detailed in Section˜3). We train the Llama 3.1 8B Instruct model for three epochs with a learning rate of 1e-6.
@@
Based on these findings, we select a batch size of 16 for our experiments in Sections˜3 and 4.
@@ B.6 Details of Evaluation
We provide the performance breakdown in terms of the category for MMLU in Table˜13.
@@ D.1 Model Selection
Currently, there are two kinds of model selections for table instruction tuning, instruction-tuning the base version of the model, as seen in works like TableLLaMA(Zhang et al., 2024a) and TableBenchLLM(Wu et al., 2024), or continuing instruction-tuning an already instruction-tuned version, as done with TableLLM(Zhang et al., 2024b) as listed in Table˜2.
As the end user may come up with their own set of instructions, we expect table instruction-tuned models to possess a strong general instruction-following ability. Imparting general instruction-following ability through table instruction-tuning to the base model is challenging, as there is a lack of diversity in the table instruction-tuning data. For instance, TableLLaMA employs six specific instruction templates across two million data points, which pales in comparison to the diverse instruction datasets in broader instruction tuning efforts such as those by Chung et al. (2024), which include 1,836 tasks, each with a set of instruction templates. As shown in Figure˜4(c), when tuning the base version of the LLaMA 3.1 8B model on instruction pairs on FeTaQA, HiTab, and TabFact, the instruction following ability of the model does not improve significantly. Moreover, with a large learning rate such as 1.0e-5, the model’s instruction following ability drops significantly when there is more training data coming in.
@@
In terms of specific table understanding tasks, tuning LLaMA 3.1 8B Instruct model yields better performance than its base version on TabFact (73.10 in Figure˜2(a) v.s. 71.10 in Figure˜4(a)) under the same experimental setup. Therefore, we select the LLaMA 3.1 8B Instruct model as our starting model.
@@ D.2 Effects of Epochs
Figure˜5 illustrates the relationship between the performance of LLaMA 3.1 8B Instruct model and the number of epochs when we fine-tune the model on the 1,500 instruction pairs at a learning rate of 1.0e-6. The model demonstrates a decent performance on these table tasks within just one or two epochs. In the meantime, the model mostly preserves its performance on MMLU and IFEval, indicating that its general capabilities are not compromised too much while acquiring table reasoning ability. Beyond this point, there is no significant performance improvement, suggesting that extending training for more epochs yields diminishing returns or may even lead to overfitting. Therefore, we choose to train our model for two epochs in Section˜4 instead of the commonly adopted six epochs by existing table LLMs as seen in Table˜2.
@@ D.3 Effects of Multi-Task
In Figure˜6, we present the heatmap of model performance when fine-tuning the LLaMA 3.1 8B Instruct model on a single dataset (one of the datasets among FeTaQA, HiTab, and TabFact). We fine-tune the model for two epochs at a learning rate of 1.0e-6 with 500 instruction pairs, and then test it against the six datasets. Additionally, Figures˜7 and 8 in Section˜D.6 present heatmaps across varying learning rates (from 1.0e-7 to 1.0e-5) and number of epochs (from one to six).
@@
There are synergy effects on these tasks. The model achieves better performance when trained on the instruction pairs combined from all three datasets, compared to being trained on each of them separately. For instance, the accuracy on HiTab increases to 66.29, compared to 64.84 when trained only on HiTab as shown in Figure˜7.
There are inter-connections between different tasks. In Figure˜6, we note that fine-tuning solely on HiTab leads to a performance of 67.80 on TabFact, and fine-tuning solely on TabFact leads to a performance of 55.62 on HiTab, demonstrating a transfer of learned capabilities between these two tasks. However, this relationship is not universal as training on HiTab yields poor performance on FeTaQA, indicating that the overlap between certain tasks may be limited.
Based on these observations, we choose to fine-tune our model on a diverse range of tasks and datasets in Section˜4.
@@ D.4 Hyperparameter Exploration Across Models
We train each model on 500 examples from HiTab, FeTaQA, and TabFact (1,500 examples total) to explore the effects of the learning rate. Table˜14 presents our results.
@@
Table˜5 lists the learning rate we would suggest for practitioners to use if they would fine-tune the LLMs on table-specific tasks.
@@
We further experiment with various training sizes for each model to observe its impact on performance. Table˜15 reports the results for Llama 2 7B, QWen 2.5, Mistral v0.3, and Phi 3 8K models at one of the learning rates we select based on our results in Table˜14.
@@
In addition, we find that given the same number of training instances, Llama 3.1 8B Instruct achieves better performance than Llama 2 7B Instruct. For instance, when trained with the same 1,500 examples at the learning rate of 1.0e-6, Llama 3.1 8B Instruct yields 73.10 on TabFact (Section˜3) while Llama 2 7B Instruct only yields 53.80 (Table˜15). Therefore, models with stronger general capabilities require less tuning data in our fine-tuning process.
@@ D.5 Hyperparameter Exploration for LoRA and QLoRA
We replicate the experiments we conduct in Section˜D.4, and here we present our results in two aspects, the learning rate and the number of examples.
@@
Table˜16 presents the results. We find that there is still a “breakdown point” where further increasing the learning rate causes a sharp decline in overall performance for both LoRA and QLoRA. However, such “breakdown point” for LoRA and QLoRA (around 5.0e-5) is larger than the full parameter tuning (usually around 1.0e-6). When the learning rate does not surpass such a “breakdown point”, both methods demonstrate competitive in-domain performance on table tasks.
@@
Table˜17 presents the results. Similar to what we have found for full parameter fine-tuning, both LoRA and QLoRA show diminishing returns as the number of training examples increases. While performance improves with more examples, the rate of improvement slows beyond 600 examples for LoRA. For QLoRA, the rate of improvement slows beyond 90 examples. We find that with 1,500 examples, QLoRA and LoRA perform similarly on the in-domain table tasks, and on FeTaQA, QLoRA even outperforms LoRA by 1 point. This suggests that practitioners may leverage such parameter-efficient fine-tuning methods like QLoRA in practice, especially when they have limited table data.
@@ D.6 Individual Task’s Influence on Model Performance
Figures˜7 and 8 present heatmaps across varying learning rates (from 1.0e-7 to 1.0e-5) and epochs (from one to six). We can see that the patterns coincide with what we have discussed in Section˜3, that a learning rate that is too large such as 1.0e-5 or too small such as 1.0e-7 leads to suboptimal table understanding ability, and the large learning rate also compromises the model’s general capabilities. Moreover, we do not observe significant performance gain when we fine-tune the model for more epochs. Across these hyperparameters, we can observe the inter-connections between tasks such as HiTab and TabFact, as training solely on one often leads to good performance on the other. But this is not universally true, as tasks such as FeTaQA and FEVEROUS seem to not have strong inter-connections.
In addition, we observe that the learning rate works the best for an individual task does not necessarily work the best for other tasks. For instance, in Figures˜7 and 8, the learning rate of 5.0e-6 yields the best performance for FeTaQA, but is suboptimal for HiTab and TabFact. This highlights that when multiple tasks are involved in the training process, researchers need to consider beyond a single task to decide their hyperparameters.
@@ D.7 Trade-off Analysis for Data Properties
Table˜18 presents the results. We find that the performance degradation is most significant on TabFact. Interestingly, despite TabFact having intermediate numeric density and table-to-question token ratios, it still shows the fastest performance decline.
@@ Appendix E Hindsight Analysis
Figure 9: Performance scores (y-axis) with respect to the number of epochs (x-axis) and learning rates. In Figure˜9(a), we aggregate the performance scores for all the datasets listed in Table˜6.
Figure˜9 provides the complete results of the model performance versus the learning rate and the number of epochs.
Apart from what we have discussed in Section˜4.3, we find that on S1, the learning rate of 5.0e-7 yields a consistent good accuracy scores (around 64 to 65) across all the epochs, while 1.0e-6 maintains a good accuracy score (around 64 to 65) for the first two epochs, but starting from the third epoch, it experiences a performance decline (from 64.93 to 52.56).
@@ Appendix F Model Prediction Examples
Table˜19 provides an example for table LLMs’ generation on IFEval dataset. Tables˜20 and 21 provide two examples for table LLM’s generation on Table-Syn dataset. Apart from the limited out-of-domain table reasoning ability, we find that existing table LLMs also exhibit limited instruction-following capabilities, and often struggle with consistently returning answers in specified formats, such as JSON. Such a limitation poses challenges in the practical use cases, where the end-users may request specific output formats to extract answers from the model’s predictions.